# A Survey on Neural Question Generation: Methods, Applications, and Prospects

### v1

@@ Abstract
In this survey, we present a detailed examination of the advancements in Neural Question Generation (NQG), a field leveraging neural network techniques to generate relevant questions from diverse inputs like knowledge bases, texts, and images. The survey begins with an overview of NQG’s background, encompassing the task’s problem formulation, prevalent benchmark datasets, established evaluation metrics, and notable applications. It then methodically classifies NQG approaches into three predominant categories: structured NQG, which utilizes organized data sources, unstructured NQG, focusing on more loosely structured inputs like texts or visual content, and hybrid NQG, drawing on diverse input modalities. This classification is followed by an in-depth analysis of the distinct neural network models tailored for each category, discussing their inherent strengths and potential limitations. The survey culminates with a forward-looking perspective on the trajectory of NQG, identifying emergent research trends and prospective developmental paths. Accompanying this survey is a curated collection of related research papers, datasets and codes, systematically organized on Github [^1] https://github.com/PersistenceForever/Neural-Question-Generation-Survey-List, providing an extensive reference for those delving into NQG.
@@ 1 Introduction
for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center, calign=edge midpoint, , [NQG, text width=2cm, for tree=fill=hidden-grey!20 [Structured NQG (§3), text width=1.8cm, for tree=fill=hidden-green!20 [Traditional Seq2Seq Models, text width=2.8cm, for tree=fill=hidden-green!20 [RNN-based, text width=2.8cm, for tree=fill=hidden-green!20 [K2Q-RNN Reddy et al. (2017); Zero-ShotQG Elsahar et al. (2018); KTG Bi et al. (2020), text width=6.25cm, node options=align=left ] ] [Transformer-based, text width=2.8cm, for tree=fill=hidden-green!20 [IDCQG Liu et al. (2019); MHQG Kumar et al. (2019a), text width=6.25cm, node options=align=left ] ] ] [Graph2Seq Models, text width=2.8cm, for tree=fill=hidden-green!20 [KGPT Chen et al. (2020a); Meta-CQG Zhang et al. (2022a); G2S Chen et al. (2023b), text width=9.5cm, node options=align=left ] ] [Pre-trained Seq2Seq Models, text width=2.8cm, for tree=fill=hidden-green!20 [PLMs-based, text width=2.8cm, for tree=fill=hidden-green!20 [JointGT Ke et al. (2021); DSM Guo et al. (2022); LFKQG Fei et al. (2022b); AutoQGS Xiong et al. (2022); DiversifyQG Guo et al. (2023), text width=6.25cm, node options=align=left ] ] [LLMs-based, text width=2.8cm, for tree=fill=hidden-green!20 [KQG-COT Liang et al. (2023), text width=6.25cm, node options=align=left ] ] ] ] [Unstructured NQG (§4), text width=1.8cm, for tree=fill=hidden-yellow!20 [Traditional Seq2Seq Models, text width=2.8cm, for tree=fill=hidden-yellow!20 [RNN-based, text width=2.8cm, for tree=fill=hidden-yellow!20 [NQGRC Du et al. (2017); HNQG Sun et al. (2018); CorefNQG Du and Claire (2018); KPEQG Subramanian et al. (2018) QGWPC Li et al. (2019); ASs2s Kim et al. (2019); CGCQG Tuan et al. (2020), text width=6.25cm, node options=align=left ] ] [Transformer-based, text width=2.8cm, for tree=fill=hidden-yellow!20 [CLQG Kumar et al. (2019b); NQGAP Wang et al. (2020a); SQG Chai and Wan (2020), text width=6.25cm, node options=align=left ] ] ] [Graph-based Models, text width=2.8cm, for tree=fill=hidden-yellow!20 [G2S-RL Chen et al. (2020b); SGGDQ Pan et al. (2020); MulQG Su et al. (2020); IGND Fei et al. (2021); CQG Fei et al. (2022a), text width=9.5cm, node options=align=left ] ] [Pre-trained Seq2Seq Models, text width=2.8cm, for tree=fill=hidden-yellow!20 [UNILM Dong et al. (2019); MiniLM Wang et al. (2020b); ERNIE Xiao et al. (2020); BertGen Back et al. (2021); MultiFactor Xia et al. (2023), text width=9.5cm, node options=align=left ] ] [Visual2Seq Models, text width=2.8cm, for tree=fill=hidden-yellow!20 [CNN-based, text width=2.8cm, for tree=fill=hidden-yellow!20 [iQAN Li et al. (2018); iVQA Liu et al. (2018); DGN Chen et al. (2021); KECVQG Chen et al. (2023a), text width=6.25cm, node options=align=left ] ] [GNN-based, text width=2.8cm, for tree=fill=hidden-yellow!20 [Radial-GCN Xu et al. (2020); MOAG Xie et al. (2021); KVQG Xie et al. (2022); MCMI Chai et al. (2023), text width=6.25cm, node options=align=left ] ] [Generative Networks-based, text width=2.8cm, for tree=fill=hidden-yellow!20 [IMVQG Krishna et al. (2019); C3VQG Uppal et al. (2020), text width=6.25cm, node options=align=left ] ] ] ] [Hybrid NQG (§5), text width=1.8cm, for tree=fill=hidden-blue!20 [Pre-trained Models, text width=2.8cm, for tree=fill=hidden-blue!20 [MultiQG-TI Wang and Baraniuk (2023) (Under-explored), text width=9.5cm, node options=align=left ] ] ] ]
@@
Figure 1: The taxonomy of NQG. We classify NQG into three types based on input modalities: structured, unstructured, and hybrid NQG.
@@ 2.2 Popular Datasets
We summarize popular datasets in NQG tasks, as shown in Table 1. Since QG can be viewed as a dual task of question answering, QG datasets are typically derived from QA datasets. To enhance clarity, we classify these datasets based on their input types, encompassing those derived from knowledge bases, natural language text, and visual sources.
Knowledge Base-based Datasets. We show three popular datasets for knowledge base question generation (KBQG). Firstly, WebQuestions (WQ) Kumar et al. (2019a) comprises 22,989 instances from WebQuestionsSP and ComplexWebQuestions, both of which are benchmarks for knowledge base question answering (KBQA). These benchmarks contain questions, answers, and corresponding SPARQL queries. Following  Kumar et al. (2019a), they convert a SPARQL query to a subgraph. Consequently, each instance in the WQ dataset includes subgraphs, answers, and questions. In addition, PathQuestions (PQ) Zhou et al. (2018) is constructed using two subsets of Freebase. Notably, in PQ, the KB subgraph forms a path between the topic entities and answer entities, typically spanning connections of 2-hop or 3-hop. Furthermore, GrailQA (GQ) Gu et al. (2021) is a large-scale, high-quality KBQA dataset. Each question is associated with an S-expression, which can be interpreted as a logical form.
@@ 2.3 Evaluation
  • BLEU. BiLingual Evaluation Understudy (BLEU) metric evaluates the average n-gram precision against the reference text, applying a penalty for excessively short text. BLEU-$n$ calculates the proportion of the common n-grams between the generated question and the ground-truth.
  • ROUGE. Recall-Oriented Understudy for Gisting Evaluation (ROUGE) focuses on recall, measuring the ratio of n-grams from the ground-truth question that are also present in the generated question.
  • METEOR. Metric for Evaluation of Translation with Explicit ORdering (METEOR) offers a more comprehensive assessment than BLEU. METEOR is calculated based on the harmonic mean of the unigram precision and recall, providing a balanced evaluation.
Diversity Metrics. The diversity metric is essential for tasks involving diversified question generation. A widely recognized metric, Distinct-$n$, calculates the proportion of unique n-grams in the generated question.
Semantic Similarity Metrics. Beyond word-level comparison, it is vital to assess sentence-level comparison, particularly about semantic similarity. BERTScore, a widely-used metric, utilizes pre-trained contextual embeddings from BERT to compare words between generated and ground-truth questions, calculating their similarity using cosine similarity.
@@ 3.0.3 Pre-trained Seq2Seq Models
PLMs-based. A significant challenge in adapting these PLMs to KBQG tasks lies in bridging the semantic gap. This is primarily because PLMs are originally pre-trained on unstructured text, which contrasts with the structured nature of KB. Additionally, another key challenge involves effectively capturing the structural information inherent in KB.
@@
To address the above challenges,  Ke et al. Ke et al. (2021) introduce three pre-training tasks including graph-enhanced text reconstruction, text-enhanced graph reconstruction, and graph-text embedding alignment to explicitly build the connection between knowledge graphs and text sequences. Additionally, they develop a structure-aware semantic aggregation module at each Transformer layer to aggregate contextual information following the graph structure.  Guo et al. Guo et al. (2022) address the issue of the semantic gap by converting the input subgraph into a linear sequence of triples, achieved through the concatenation of relational triples. Furthermore, they explicitly incorporate structural information as input. This incorporation involves directly inserting special tokens “⟨H⟩”, “⟨R⟩”, and “⟨T⟩” before the head entity, relation, and tail entity in each fact triple, respectively. This approach effectively clarifies the relationships between entities, ensuring a more coherent representation within the model. Likewise,  Guo et al. Guo et al. (2023) add special tokens indicative of structured information at the beginning of each element within every triple in KB.  Xiong et al. Xiong et al. (2022) focus on directly generating questions from SPARQL, aimed at covering complex operations. They first execute the SPARQL query on a KB to retrieve a corresponding subgraph. This subgraph is then linearized, serving as input for an auto-prompter, which generates the prompt text. Subsequently, this prompt text, along with the original SPARQL query, are used as inputs for a QG generator. This process ultimately results in a high-quality question generation, effectively bridging the gap between non-natural language SPARQL and the natural language question.
@@
LLMs-based. Despite the success of PLMs-based models on KBQG Fei et al. (2022b); Guo et al. (2022), their effectiveness hinges on extensive fine-tuning using large training datasets. However, the creation of labeled datasets is costly and time-consuming. Hence, researchers are increasingly focusing on few-shot KBQG tasks to mitigate these challenges Xiong et al. (2022). Recently, LLMs, like ChatGPT and Llama2, have exhibited exceptional capabilities in various few-shot and zero-shot tasks. This emerging insight inspires researchers to investigate few-shot KBQG tasks, leveraging the capabilities of LLMs. The key challenge is to design effective prompts that prompt LLMs to generate the targeted questions for KBQG.  Liang et al. (2023) propose KQG-COT framework, which involves first employing LLMs (i.e., text-davinci-003) to generate ideal questions for KBQG using chain-of-thought (COT). To be specific, KQG-COT first identifies suitable logical forms from the unlabeled data pool, meticulously evaluating their attributes. Following this, a specialized prompt is developed to steer LLMs in creating complex questions derived from these chosen logical forms.
@@ 4.1.1 Traditional Seq2Seq Models
RNN-based. Du et al. Du et al. (2017) first apply RNN to TQG tasks, leveraging an attention mechanism to enable the decoder to concentrate on the most pertinent segments of the input text. To decide which information to focus on when generating questions, most Seq2Seq models leverage the answer position features to incorporate the answer spans. For example,  Sun et al. Sun et al. (2018) contend that context words near the answer are more apt to be answer-relevant. Hence, they explicitly encode the positional proximity of these context words to the answer by position embedding and a position-aware attention mechanism. Nevertheless,  Li et al. Li et al. (2019) think the proximity-based approach does not always work. Therefore, they devise a more generalized model, which exploits answer-relevant relations to facilitate the faithfulness of the generated question.
@@ 4.1.3 Pre-trained Seq2Seq Models
To address these challenges, researchers are increasingly focusing on fine-tuning PLMs for specific downstream tasks and adapting their neural architectures accordingly Dong et al. (2019); Wang et al. (2020b); Xiao et al. (2020); Xia et al. (2023). For example,  Dong et al. Dong et al. (2019) present a unified pre-trained language model (UNILM), which is distinctively optimized across three distinct types of language modeling tasks, including unidirectional, bidirectional, and sequence-to-sequence prediction.  Back et al. Back et al. (2021) propose a novel pre-training approach tailored specifically for QG tasks. This approach intensively focuses on the answer, aiming to generate contextually relevant sentences containing missing answers. By doing this, it aims to learn more effective representations that are highly optimized for the question generation task.  Xia et al. Xia et al. (2023) introduce phrase-enhanced Transformer, an effective model that capitalizes on the strengths of powerful PLMs. This method creatively integrates phrase selection probabilities from the encoder into the decoder, significantly enhancing the quality of question generation.

### v2

@@ Abstract
In this survey, we present a detailed examination of the advancements in Neural Question Generation (NQG), a field leveraging neural network techniques to generate relevant questions from diverse inputs like knowledge bases, texts, and images. The survey begins with an overview of NQG’s background, encompassing the task’s problem formulation, prevalent benchmark datasets, established evaluation metrics, and notable applications. It then methodically classifies NQG approaches into three predominant categories: structured NQG, which utilizes organized data sources, unstructured NQG, focusing on more loosely structured inputs like texts or visual content, and hybrid NQG, drawing on diverse input modalities. This classification is followed by an in-depth analysis of the distinct neural network models tailored for each category, discussing their inherent strengths and potential limitations. The survey culminates with a forward-looking perspective on the trajectory of NQG, identifying emergent research trends and prospective developmental paths. Accompanying this survey is a curated collection of related research papers, datasets and codes [^1] https://github.com/PersistenceForever/Neural-Question-Generation-Survey-List, providing an extensive reference for those delving into NQG.
@@ 1 Introduction
for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center, calign=edge midpoint, , [NQG, text width=2cm, for tree=fill=hidden-grey!20 [Structured NQG (§3), text width=1.8cm, for tree=fill=hidden-green!20 [Traditional Seq2Seq Models, text width=2.8cm, for tree=fill=hidden-green!20 [RNN-based, text width=2.8cm, for tree=fill=hidden-green!20 [K2Q-RNN Reddy et al. (2017); Zero-ShotQG Elsahar et al. (2018); KTG Bi et al. (2020), text width=6.25cm, node options=align=left ] ] [Transformer-based, text width=2.8cm, for tree=fill=hidden-green!20 [IDCQG Liu et al. (2019); MHQG Kumar et al. (2019a), text width=6.25cm, node options=align=left ] ] ] [Graph2Seq Models, text width=2.8cm, for tree=fill=hidden-green!20 [KGPT Chen et al. (2020a); Meta-CQG Zhang et al. (2022a); G2S Chen et al. (2023b), text width=9.5cm, node options=align=left ] ] [Pre-trained Seq2Seq Models, text width=2.8cm, for tree=fill=hidden-green!20 [PLMs-based, text width=2.8cm, for tree=fill=hidden-green!20 [JointGT Ke et al. (2021); DSM Guo et al. (2022); LFKQG Fei et al. (2022b); AutoQGS Xiong et al. (2022); DiversifyQG Guo et al. (2023), text width=6.25cm, node options=align=left ] ] [LLMs-based, text width=2.8cm, for tree=fill=hidden-green!20 [KQG-COT Liang et al. (2023); SGSH Guo et al. (2024), text width=6.25cm, node options=align=left ] ] ] ] [Unstructured NQG (§4), text width=1.8cm, for tree=fill=hidden-yellow!20 [Traditional Seq2Seq Models, text width=2.8cm, for tree=fill=hidden-yellow!20 [RNN-based, text width=2.8cm, for tree=fill=hidden-yellow!20 [NQGRC Du et al. (2017); HNQG Sun et al. (2018); CorefNQG Du and Claire (2018); KPEQG Subramanian et al. (2018) QGWPC Li et al. (2019); ASs2s Kim et al. (2019); CGCQG Tuan et al. (2020), text width=6.25cm, node options=align=left ] ] [Transformer-based, text width=2.8cm, for tree=fill=hidden-yellow!20 [CLQG Kumar et al. (2019b); NQGAP Wang et al. (2020a); SQG Chai and Wan (2020), text width=6.25cm, node options=align=left ] ] ] [Graph-based Models, text width=2.8cm, for tree=fill=hidden-yellow!20 [G2S-RL Chen et al. (2020b); SGGDQ Pan et al. (2020); MulQG Su et al. (2020); IGND Fei et al. (2021); CQG Fei et al. (2022a), text width=9.5cm, node options=align=left ] ] [Pre-trained Seq2Seq Models, text width=2.8cm, for tree=fill=hidden-yellow!20 [UNILM Dong et al. (2019); MiniLM Wang et al. (2020b); ERNIE Xiao et al. (2020); BertGen Back et al. (2021); CoHS-CQG Do et al. (2022); mQG Yoon and Bak (2023); SG-CQG Do et al. (2023); MultiFactor Xia et al. (2023), text width=9.5cm, node options=align=left ] ] [Visual2Seq Models, text width=2.8cm, for tree=fill=hidden-yellow!20 [CNN-based, text width=2.8cm, for tree=fill=hidden-yellow!20 [iQAN Li et al. (2018); iVQA Liu et al. (2018); DGN Chen et al. (2021); KECVQG Chen et al. (2023a), text width=6.25cm, node options=align=left ] ] [GNN-based, text width=2.8cm, for tree=fill=hidden-yellow!20 [Radial-GCN Xu et al. (2020); MOAG Xie et al. (2021); KVQG Xie et al. (2022); MCMI Chai et al. (2023), text width=6.25cm, node options=align=left ] ] [Generative Networks-based, text width=2.8cm, for tree=fill=hidden-yellow!20 [IMVQG Krishna et al. (2019); C3VQG Uppal et al. (2020), text width=6.25cm, node options=align=left ] ] ] ] [Hybrid NQG (§5), text width=1.8cm, for tree=fill=hidden-blue!20 [MultiQG-TI Wang and Baraniuk (2023); ConVQG Mi et al. (2024); UniCFQG Dong et al. (2024) (Under-explored), text width=12.75cm, node options=align=left ] ] ]
@@
Figure 1: The taxonomy of NQG. We classify NQG into three types based on input modalities: Structured NQG, which deals with structured data; Unstructured NQG, which handles unstructured data; and Hybrid NQG, which integrates both structured and unstructured data.
@@ 2.2 Popular Datasets
We summarize popular datasets in NQG tasks, as shown in Table 1. Since QG can be viewed as a dual task of QA, QG datasets are typically derived from QA datasets. To enhance clarity, we classify these datasets based on their input types, encompassing those derived from knowledge bases, natural language text, and visual sources.
Knowledge Base-based Datasets. We show three popular datasets for knowledge base question generation (KBQG). Firstly, WebQuestions (WQ) Kumar et al. (2019a) comprises 22,989 instances from WebQuestionsSP Yih et al. (2016) and ComplexWebQuestions Talmor and Berant (2018), both of which are benchmarks for knowledge base question answering (KBQA). These benchmarks contain questions, answers, and corresponding SPARQL queries. Following  Kumar et al. (2019a), they convert a SPARQL query to a subgraph. Consequently, each instance in the WQ dataset includes subgraphs, answers, and questions. In addition, PathQuestions (PQ) Zhou et al. (2018) is constructed using two subsets of Freebase. Notably, in PQ, the KB subgraph forms a path between the topic entities and answer entities, typically spanning connections of 2-hop or 3-hop. Furthermore, GrailQA (GQ) Gu et al. (2021) is a large-scale, high-quality KBQA dataset. Each question is associated with an S-expression, which can be interpreted as a logical form.
@@ 2.3 Evaluation
  • BLEU. BiLingual Evaluation Understudy (BLEU) Papineni et al. (2002) metric evaluates the average n-gram precision against the reference text, applying a penalty for excessively short text. BLEU-$n$ calculates the proportion of the common n-grams between the generated question and the ground-truth.
  • ROUGE. Recall-Oriented Understudy for Gisting Evaluation (ROUGE) Lin (2004) focuses on recall, measuring the ratio of n-grams from the ground-truth question that are also present in the generated question.
  • METEOR. Metric for Evaluation of Translation with Explicit ORdering (METEOR) Banerjee and Lavie (2005) offers a more comprehensive assessment than BLEU. METEOR is calculated based on the harmonic mean of the unigram precision and recall, providing a balanced evaluation.
Diversity Metrics. The diversity metric is essential for tasks involving diversified question generation. A widely recognized metric, Distinct-$n$ Li et al. (2016), calculates the proportion of unique n-grams in the generated question.
Semantic Similarity Metrics. Beyond word-level comparison, it is vital to assess sentence-level comparison, particularly about semantic similarity. BERTScore Zhang et al. (2020), a widely-used metric, utilizes pre-trained contextual embeddings from BERT to compare words between generated and ground-truth questions, calculating their similarity using cosine similarity.
@@ 3.0.3 Pre-trained Seq2Seq Models
PLMs-based. BART Lewis et al. (2020) and T5 Raffel et al. (2020), two widely recognized PLMs, are built upon the encoder-decoder framework. A significant challenge in adapting these PLMs to KBQG tasks lies in bridging the semantic gap. This is primarily because PLMs are originally pre-trained on unstructured text, which contrasts with the structured nature of KB. Additionally, another key challenge involves effectively capturing the structural information inherent in KB.
@@
To address the above challenges,  Ke et al. Ke et al. (2021) introduce three pre-training tasks: graph-enhanced text reconstruction, text-enhanced graph reconstruction, and graph-text embedding alignment to explicitly build the connection between knowledge graphs and text sequences. Additionally, they create a semantic aggregation module that is aware of the structure at every Transformer layer, which gathers contextual information according to the structure of the graph.  Guo et al. Guo et al. (2022) address the issue of the semantic gap by converting the input subgraph into a linear sequence of triples, achieved through the concatenation of relational triples. Furthermore, they explicitly incorporate structural information as input. This incorporation involves directly inserting special tokens “⟨H⟩”, “⟨R⟩”, and “⟨T⟩” before the head entity, relation, and tail entity in each fact triple, respectively. This approach effectively clarifies the relationships between entities, ensuring a more coherent representation within the model. Likewise,  Guo et al. Guo et al. (2023) add special tokens indicative of structured information at the beginning of each element within every triple in KB.  Xiong et al. Xiong et al. (2022) focus on directly generating questions from SPARQL, aimed at covering complex operations. They first execute the SPARQL query on a KB to retrieve a corresponding subgraph. This subgraph is then linearized, serving as input for an auto-prompter, which generates the prompt text. Subsequently, this prompt text, along with the original SPARQL query, are used as inputs for a QG generator. This process ultimately results in high-quality question generation, effectively bridging the gap between non-natural language SPARQL and the natural language question.
@@
LLMs-based. Despite the success of PLMs-based models on KBQG Fei et al. (2022b); Guo et al. (2022, 2023), their effectiveness hinges on extensive fine-tuning using large training datasets. However, the creation of labeled datasets is costly and time-consuming. Hence, researchers are increasingly focusing on few-shot KBQG tasks to mitigate these challenges Xiong et al. (2022). Recently, LLMs, like ChatGPT and Llama2, have exhibited exceptional capabilities in various few-shot and zero-shot tasks. This emerging insight inspires researchers to investigate few-shot KBQG tasks, leveraging the capabilities of LLMs. The key challenge is to design effective prompts that prompt LLMs to generate the targeted questions for KBQG.  Liang et al. Liang et al. (2023) propose KQG-COT framework, which involves first employing LLMs (i.e., text-davinci-003) to generate ideal questions for KBQG using chain-of-thought (COT). To be specific, KQG-COT first identifies suitable logical forms from the unlabeled data pool, meticulously evaluating their attributes. Following this, a specialized prompt is developed to steer LLMs in creating complex questions derived from these chosen logical forms.  Guo et al. Guo et al. (2024) develop a fine-grained prompting approach named SGSH. Concretely, SGSH involves training a learnable skeleton generator that then uses the generated skeleton to create skeleton-based prompts, effectively stimulating LLMs to generate desired questions.
@@ 4.1.1 Traditional Seq2Seq Models
RNN-based. Du et al. Du et al. (2017) first apply RNN to TQG tasks, leveraging an attention mechanism to enable the decoder to concentrate on the most pertinent segments of the input text. In the process of determining the specific information to emphasize while generating questions, the majority of Seq2Seq models utilize the features of answer positions to integrate the spans of answers. For example,  Sun et al. Sun et al. (2018) contend that context words near the answer are more apt to be answer-relevant. Hence, they explicitly encode the positional proximity of these context words to the answer by position embedding and a position-aware attention mechanism. Nevertheless,  Li et al. Li et al. (2019) think the proximity-based approach does not always work. Therefore, they devise a more generalized model, which exploits answer-relevant relations to facilitate the faithfulness of the generated question.
@@ 4.1.3 Pre-trained Seq2Seq Models
To address these challenges, researchers are increasingly focusing on fine-tuning PLMs for specific downstream tasks and adapting their neural architectures accordingly Dong et al. (2019); Wang et al. (2020b); Xiao et al. (2020); Do et al. (2022); Xia et al. (2023). For example,  Dong et al. Dong et al. (2019) present a unified pre-trained language model (UNILM), which is distinctively optimized across three distinct types of language modeling tasks, including unidirectional, bidirectional, and sequence-to-sequence prediction.  Back et al. Back et al. (2021) propose a novel pre-training approach tailored specifically for QG tasks. This approach intensively focuses on the answer, aiming to generate contextually relevant sentences containing missing answers. By doing this, it aims to learn more effective representations that are highly optimized for the question generation task.  Do et al. Do et al. (2023) present a new framework comprising two modules for generating conversational questions: “what-to-ask” and “how-to-ask”. The “what-to-ask” module constructs a semantic graph to extract underlying rationale and selects the relevant answer span. The “how-to-ask” module uses a classifier to identify the appropriate question type. Subsequently, the framework fine-tunes the T5 Raffel et al. (2020) model on the tailored dataset to produce conversational questions.  Xia et al. Xia et al. (2023) introduce phrase-enhanced Transformer, an effective model that capitalizes on the strengths of powerful PLMs. This method creatively integrates phrase selection probabilities from the encoder into the decoder, significantly enhancing the quality of question generation.